OpenStudy (anonymous):

solve algebraically: x-y+z=5 -2x-y-z=-6 3x+3y-2z=-5 please and thank you :)

OpenStudy (anonymous):

Well, first you need to figure out one variable: x-y+z=5 -2x-y-z=-6 If you subtract the second from the first what do you receive?

OpenStudy (anonymous):

-3x-2y-2z=-11?

OpenStudy (anonymous):

I tried this once before but it didn't work out right so now I'm unsure about my methods/math

OpenStudy (anonymous):

You need to subtract them, you are adding them together.

OpenStudy (anonymous):

multiply both side of 3x + 3y - 2z = -5 by -1 -3x -3y + 2z = 5 set it equal to x-y+z = 5 -3x -3y + 2z = x - y + z 4x + 2y -z = 0 for this equation -2x - y -z =-6 add 6 on both sides -2x - y - z + 6 = 0 set the 2 equation equal to each other 4x + 2y -z = -2x - y - z +6 simplify 6x + 3y = 6 solve for x 6x = 6 -3y x = 1 - y/3 substitute back in the equation 6x + 3y = 6 6(1 - y/3) + 3y = 6 6 - 2y + 3y = 6 y = 0 that means x = 1 and x - y + z = 5 1 - 0 + z = 5 z = 4

OpenStudy (anonymous):

Thank you so much! I was so stuck on that one! And thank you as well malical for your time :)

OpenStudy (anonymous):

np glad to help